OpenStudy (bethb):

The general form of the equation of a circle is x2 + y2 + 42x + 38y − 47 = 0. The equation of this circle in standard form is....... The center of the circle is at the point........ and its radius is...... units.

OpenStudy (jango_in_dtown):

complete the squares.

OpenStudy (jango_in_dtown):

\[x^2+y^2+42x+38y-47=0 \implies (x+21)^2+(y+19)^2=47+21^2+19^2\]

OpenStudy (jango_in_dtown):

centre is ???

OpenStudy (welshfella):

can you find where the centre is from the equation?

OpenStudy (bethb):

sorry that was homework for the other day